# Asymptotic Mean of Digits of the $Q_{s}$–representation of the Fractional Part of a Real Number and Related Problems of Fractal Geometry and Fractal Analysis

### 03 Mar 2026 v1

@@ 1. Introduction
A number with a periodic $Q_{s}$-representation has a well-defined asymptotic mean its digits. For example, $r\left(\Delta^{Q_{s}}_{(i)}\right)=i$, $r\left(\Delta^{Q_{s}}_{(ij)}\right)=\dfrac{i+j}{2}$, where $i,j\in\mathcal{A}_{s}$.
It follows from the definition that the function $r$ takes values in the interval [0,s1][0,s-1]. We note that the value of $r(x)$ does not depend on any finite number of $Q_{s}$–symbols of $x$.
@@
$x_{0}=\Delta^{Q_{s}}_{dcddcc\ldots\underbrace{d\ldots d}_{2^{k-1}}\underbrace{c\ldots c}_{2^{k-1}}\ldots},\,\,\,\,\text{де}\,\,\,\,c,d\in\mathcal{A}_{s}.$
Thus, the asymptotic mean function $r$ is defined on an everywhere dense subset of the interval $[0,1]$. On another everywhere dense subset of the interval $[0,1]$, it is undefined.
@@
Let $N_{i}(x,k)$ be the number of occurrences of the digit $i\in\mathcal{A}_{s}$ in the $Q_{s}$–representation $\Delta^{Q_{s}}_{\alpha_{1}\alpha_{2}\ldots\alpha_{k}\ldots}$ of a number $x\in[0,1]$ up to and including the kk–th position, that is,
@@
The frequency (asymptotic frequency) of the symbol $i$ in the $Q_{s}$–representation of a number $x$ is defined as the limit (if it exists)
@@
Clearly, if a number has period $(c_{1}c_{2}\ldots c_{m})$, with $c_{i}\in\mathcal{A}_{s}$, then each symbol has a well–defined frequency [17].
@@ 4. Normality of Numbers and the Role of Normality in the Theory of Singular Functions
A number $x$ is called normal in base $s$ (weakly normal) if for each $i\in\mathcal{A}_{s}$ the frequency exists and satisfies $\nu_{i}(x)=s^{-1}$. A number $x$ that is normal in every natural base $s\geq 2$ is called normal.
@@
The concept of a normal number was introduced by É. Borel in 1909. Borel’s classical theorem [3] asserts that The set of normal numbers (the set of numbers in the interval $[0,1]$ whose ss–adic representations have all digit frequencies equal to $s^{-1}$) has full Lebesgue measure. Thus, almost all real numbers are normal.
@@
The set $S$ is a continuous, everywhere dense, and everywhere discontinuous set of zero Lebesgue measure. The set $S$ is superfractal, that is, its Hausdorff–Besicovitch dimension $\alpha_{0}(S)$ equals $1$.
@@
Clearly, distinct pairs $x_{0}$ and $x^{\prime}_{0}$ correspond to distinct numbers $f(x_{0})$ and $f(x^{\prime}_{0})$. Since we can choose $x_{0}$ in a continuum of ways, the resulting set of numbers $\hat{x}$ is also continuous. In this way, we construct a mapping from the set $S$ onto a certain continuous set $S^{\prime}$. Hence, the set $S$ is continuous.
@@
Lebesgue Measure. The set $S$ consists of all numbers црщiу limit $\lim\limits_{n\to\infty}\frac{1}{n}\sum\limits_{i=1}^{n}\alpha_{i}(x)$ does not exist, in other words the limit
@@
Hausdorff–Besicovitch Dimension. From the conditions $\nu_{0}(x)+\nu_{1}(x)+\nu_{2}(x)=1$ and $\nu_{0}=\nu_{1}(x)+2\nu_{2}(x)$ we determine $\nu_{1}=\dfrac{1}{2}-\dfrac{3}{2}\nu_{2},\>\>\nu_{0}=\dfrac{1}{2}-\dfrac{3}{2}\nu_{2}+2\nu_{2}=\dfrac{1}{2}+\dfrac{1}{2}\nu_{2}.$
@@
Using Cardano’s formulas, we obtain
@@
Hausdorff–Besicovitch Dimension. From the conditions $\nu_{0}(x)+\nu_{1}(x)+\nu_{2}(x)=1$ i $\nu_{2}(x)=0$ we find $\nu_{1}(x)=1-\nu_{0}(x).$
@@
Clearly, the function attains its maximum at $\nu_{0}(x)=\nu_{1}(x)=\dfrac{1}{2}$. Thus,
@@
Everywhere density. The set $M_{2}$ is everywhere dense since every cylinder $\Delta^{3}_{c_{1}\ldots c_{k}}=\{x:\alpha_{j}(x)=c_{j},\>j=\overline{1,k}\}$ contains points from $M_{2}$. Specifically, if $x_{0}=\Delta^{3}_{\alpha_{1}\alpha_{2}\ldots\alpha_{n}\ldots}\in M_{1}$ then the point $x=\Delta^{3}_{c_{1}\ldots c_{k}\alpha_{1}\ldots\alpha_{n}\ldots}\in M_{1}\cap\Delta^{3}_{c_{1}\ldots c_{k}}$.

### 05 Mar 2026 v2

@@ 1. Introduction
A number with a periodic $Q_{s}$–representation has a well-defined asymptotic mean of its digits. For example, $r\left(\Delta^{Q_{s}}_{(i)}\right)=i$, $r\left(\Delta^{Q_{s}}_{(ij)}\right)=\dfrac{i+j}{2}$, where $i,j\in\mathcal{A}_{s}$.
It follows from the definition that the function $r$ takes values in the interval [0,s1][0,s-1]. We note that the value of $r(x)$ does not depend on any finite quantity of $Q_{s}$–symbols of $x$.
@@
$x_{0}=\Delta^{Q_{s}}_{dcddcc\ldots\underbrace{d\ldots d}_{2^{k-1}}\underbrace{c\ldots c}_{2^{k-1}}\ldots},\,\,\,\,\text{where}\,\,\,\,c,d\in\mathcal{A}_{s}.$
Thus, the asymptotic mean function $r$ is defined on an everywhere dense subset of the interval $[0,1]$. On another everywhere dense subset of the interval $[0,1]$ it is undefined.
@@
Let $N_{i}(x,k)$ be the quantity of the digits $i\in\mathcal{A}_{s}$ in the $Q_{s}$–representation $\Delta^{Q_{s}}_{\alpha_{1}\alpha_{2}\ldots\alpha_{k}\ldots}$ of a number $x\in[0,1]$ up to and including the kk–th position, that is,
@@
The frequency (asymptotic frequency) of the symbol $i$ in the $Q_{s}$–representation of the number $x$ is defined as the limit (if it exists)
@@
As usual, if a number has period $(c_{1}c_{2}\ldots c_{m})$, with $c_{i}\in\mathcal{A}_{s}$, then each symbol has a well–defined frequency [17].
@@ 4. Normality of numbers and the role of normality in the theory of singular functions
A number $x$ is called normal on the base $s$ (weakly normal) if for each $i\in\mathcal{A}_{s}$ the frequency exists and satisfies $\nu_{i}(x)=s^{-1}$. A number $x$ that is normal in every natural base $s\geq 2$ is called normal.
@@
The concept of a normal numbers was introduced by É. Borel in 1909. Borel’s classical theorem [3] asserts that The set of normal numbers (the set of numbers from the interval $[0,1]$ whose ss–adic representations have all digit frequencies equal to $s^{-1}$) has full Lebesgue measure. Thus, almost all real numbers are normal.
@@
The set $S$ is a continuous, everywhere dense and everywhere discontinuous set of zero Lebesgue measure. The set $S$ is superfractal, that is, its Hausdorff–Besicovitch dimension $\alpha_{0}(S)$ equals $1$.
@@
Distinct pairs $x_{0}$ and $x^{\prime}_{0}$ correspond to distinct numbers $f(x_{0})$ and $f(x^{\prime}_{0})$. Since we can choose $x_{0}$ in a continuum of ways then the resulting set of numbers $\hat{x}$ is also continuous. In this way we construct a mapping from the set $S$ onto a certain continuous set $S^{\prime}$. Hence, the set $S$ is continuous.
@@
Lebesgue Measure. The set $S$ consists of all numbers whose limit $\lim\limits_{n\to\infty}\frac{1}{n}\sum\limits_{i=1}^{n}\alpha_{i}(x)$ does not exist, in other words the limit
@@
Hausdorff–Besicovitch dimension. From the conditions $\nu_{0}(x)+\nu_{1}(x)+\nu_{2}(x)=1$ and $\nu_{0}=\nu_{1}(x)+2\nu_{2}(x)$ we determine $\nu_{1}=\dfrac{1}{2}-\dfrac{3}{2}\nu_{2}$ and $\nu_{0}=\dfrac{1}{2}-\dfrac{3}{2}\nu_{2}+2\nu_{2}=\dfrac{1}{2}+\dfrac{1}{2}\nu_{2}.$
@@
Using Cardano’s formulas we obtain
@@
Hausdorff–Besicovitch dimension. From the conditions $\nu_{0}(x)+\nu_{1}(x)+\nu_{2}(x)=1$ and $\nu_{2}(x)=0$ we find $\nu_{1}(x)=1-\nu_{0}(x).$
@@
The function attains its maximum at $\nu_{0}(x)=\nu_{1}(x)=\dfrac{1}{2}$. Thus,
@@
Everywhere density. The set $M_{2}$ is everywhere dense set since every cylinder $\Delta^{3}_{c_{1}\ldots c_{k}}=\{x:\alpha_{j}(x)=c_{j},\>j=\overline{1,k}\}$ contains points from $M_{2}$. Specifically, if $x_{0}=\Delta^{3}_{\alpha_{1}\alpha_{2}\ldots\alpha_{n}\ldots}\in M_{1}$ then the point $x=\Delta^{3}_{c_{1}\ldots c_{k}\alpha_{1}\ldots\alpha_{n}\ldots}\in M_{1}\cap\Delta^{3}_{c_{1}\ldots c_{k}}$.